# Logarithmic Enriques varieties

### v1

@@ 6.1. Known examples of Enriques manifolds
There is a wide diversity of geometric constructions of Enriques surfaces, we refer for instance to [23, 25, 44]. However, in higher dimension quite few geometric constructions of nonsingular Enriques varieties do exist. We briefly review the constructions given in [15, 47] and give some new examples using similar methods.
For any two-dimensional complex torus $A$ with origin $0\in A$, we denote by ${A}^{(n)}$ the $n$-th symmetric power of $A$, by $s\colon{A}^{(n)}\to A$ the summation map and by ${A}^{[n]}$ the Hilbert scheme of $n$ points on $A$. Consider the Hilbert–Chow morphism $\rho\colon{A}^{[n]}\to{A}^{(n)}$, the $n$-th generalized Kummer variety of $A$ is the fiber ${\rm K}_{n}(A)\coloneqq(s\circ\rho)^{-1}(0)$.
@@
Let $E$ be an Enriques surface. Then for any $n\geq 2$, the Hilbert scheme ${E}^{[n]}$ of $n$ point on $E$ is an Enriques manifold of Calabi–Yau type, of dimension $2n$ and index two [47, Theorem 3.1].
@@
Let $A=E\times F$ be the product of two elliptic curves and consider the automorphism $\iota_{A}$ of $A$ defined by $\iota_{A}(x,y)=(-x+u,y+v)$. For any integer $n\geq 1$, the natural automorphism ${\iota_{A}}^{[n+1]}$ of ${A}^{[n+1]}$ respects the fiber ${\rm K}_{n}(A)$ if $u$ and $v$ are $(n+1)$-torsion points of $E$ and $F$ respectively, and $\iota_{A}$ is an involution on $A$ if $v$ is a $2$-torsion point. Assuming thus that $n$ is odd, $u\in E[n+1]$ and $v\in F[2]$, it is easy to check that the natural automorphism ${\rm K}_{n}(\iota_{A})$ is a fixed point free involution on ${\rm K}_{n}(A)$ if $u$ is not an $\frac{n+1}{2}$-torsion point, so under these conditions, if $n$ is odd, by Propositions 5.1&5.4 the quotient ${\rm K}_{n}(A)/{\rm K}_{n}(\iota_{A})$ is an Enriques manifold of symplectic type, of dimension $2n$ and index two.
@@ 6.1.3. Enriques involutions on K3 surfaces
Let $S$ be a K3 surface carrying an Enriques involution $\iota_{S}$, i.e.a fixed point free involution, and let $n\geq 1$ be an odd integer. Consider the natural involution ${\iota_{S}}^{[n]}$ induced on the Hilbert scheme ${S}^{[n]}$. Then, when $n$ is odd, the quotient ${S}^{[n]}/{\iota_{S}}^{[n]}$ is an Enriques manifold of dimension $2n$, of symplectic type and index two.
@@
Let $S$ be a very general $K3$ surface carrying an Enriques involution $\iota_{S}$ as above By genericity we can assume that $\rho(S)=10$ and that the Picard lattice $\mathrm{Pic}(S)$ is $\iota_{S}$-invariant. Given a primitive Mukai vector $v=(r,l,\chi-r)\in H^{*}(S,\mathbb{Z})$ and a $v$-general polarization $H\in\mathrm{Pic}(S)$, the induced involution $\iota_{S}^{*}$ acts on the moduli space $M_{v,H}(S)$ of $H$-stable sheaves on $S$ with Mukai vector $v$. If $\chi$ is an odd integer, the quotient $M_{v,H}(S)/\iota_{S}^{*}$ is an Enriques manifold of dimension $v^{2}+2$, of symplectic type and index two [47, Theorem 5.3]. This is a generalization of the previous construction.
@@ 6.1.5. Complete intersections of quadrics
For any even integer $n\geq 1$, consider the projective space $\mathbb{P}^{2n+1}$ with homogeneous coordinates $[x_{0}:\ldots:x_{n}:y_{0}:\ldots:y_{n}]$. For generic quadrics $Q_{j}(x),\tilde{Q}_{j}(y)$ for $j=1,\ldots,n+1$, the complete intersection:
$Y\coloneqq\{[x:y]\,|\,Q_{j}(x)=\tilde{Q}_{j}(y),\quad\forall j=1,\ldots,n+1\}$
is a nonsingular Calabi–Yau variety. The involution $\iota$ of $\mathbb{P}^{2n+1}$ defined by $\iota(x_{i})=-x_{i}$ and $\iota(y_{i})=y_{i}$ for $i=0,\ldots,n$ fixes the two $n$-dimensional projective spaces defined by $\{x_{i}=0\,|\,i=0,\ldots,n\}$ and $\{y_{i}\,|\,i=0,\ldots,n\}$. For generic choices of the quadrics, this fixed locus does not meet $Y$ so the quotient $Y/\iota$ is an Enriques manifold of Calabi–Yau type, of dimension $n$ and index two. Observe that if $n$ is odd, the same construction gives an involution preserving the volume form so the quotient is a Calabi–Yau manifold.
@@
Let $A=E\times E_{\omega}$ where $E$ is an elliptic curve, $\omega$ is a primitive $3$rd root of unity and $E_{\omega}=\frac{\mathbb{C}}{\mathbb{Z}\oplus\omega\mathbb{Z}}$. Take $u\in E$, $v\in E_{\omega}$ and consider the automorphism $f_{A}(x,y)=(x+u,\omega y+v)$. Assume that $u$ is a $3$-torsion point, so that $f_{A}$ has order $3$. Assuming that $u$ and $v$ are $(n+1)$-torsion points, the natural automorphism ${f_{A}}^{[n+1]}$ of ${A}^{[n+1]}$ respects the fiber ${\rm K}_{n}(A)$ and the necessary condition given by Proposition 5.1 imposes that $3$ divides $n+1$, so we write $n+1=3m$. If $m(2+\omega)v\neq 0$, the natural automorphism ${\rm K}_{n}(f_{A})$ is a fixed point free involution on ${\rm K}_{n}(A)$ and by Proposition 5.4, under these assumptions, if $3$ divides $n+1$ the quotient ${\rm K}_{n}(A)/{\rm K}_{n}(f_{A})$ is an Enriques manifold of symplectic type, of dimension $2n$ and index three.
A variant of this construction, starting from a quotient $A=\frac{E\times E_{\omega}}{\mathbb{Z}/3\mathbb{Z}}$, produces for an appropriate choice of the automorphism a different family of Enriques manifolds of symplectic type as quotients of generalized Kummer manifolds, see [14, §4.3.7 - Type 6].
@@
Similarly as above, take $A=E\times E_{\mathrm{i}}$ where $E_{\mathrm{i}}=\frac{\mathbb{C}}{\mathbb{Z}\oplus\mathrm{i}\mathbb{Z}}$. Take $u\in E$, $v\in E_{\mathrm{i}}$ and consider the automorphism $f_{A}(x,y)=(x+u,\mathrm{i}y+v)$. Assume that $u$ is a $4$-torsion point, so that $f_{A}$ has order $4$. Assuming that $u$ and $v$ are $(n+1)$-torsion points, the natural automorphism ${f_{A}}^{[n+1]}$ of ${A}^{[n+1]}$ respects the fiber ${\rm K}_{n}(A)$ and the necessary condition given by Proposition 5.1 imposes that $4$ divides $n+1$, so we write $n+1=4m$. If $2mu\neq 0$ or $2m(1+\mathrm{i})v\neq 0$, the natural automorphism ${\rm K}_{n}(f_{A})$ is a fixed point free involution on ${\rm K}_{n}(A)$ and by Proposition 5.4, under these assumptions, if $4$ divides $n+1$ the quotient ${\rm K}_{n}(A)/{\rm K}_{n}(f_{A})$ is an Enriques manifold of symplectic type, of dimension $2n$ and index four.
@@ 6.2. Log-Enriques varieties of IHS type
By Proposition 5.4, any quotient of a $2n$-dimensional IHS manifold by a purely non-symplectic automorphism is a log-Enriques variety as long as the order of the automorphism does not divide $n$. Here, we discuss some features of examples which can be constructed from some deformation families of IHS manifolds.
@@ 6.2.1. Log-Enriques varieties of prime index as quotients of IHS manifolds
Nonsymplectic automorphisms of prime order acting on IHS manifolds have now been classified for all known deformation families (see [12, 11, 14, 19, 20, 17] for $K3^{[n]}$ type, [43] for generalized Kummer manifolds, [32] for $OG6$ manifolds and [17, 10] for $OG10$ manifolds). Let $Y$ be a manifold and let $\varphi\in\mathrm{Aut}(Y)$ be nonsymplectic of prime order $p$; in particular, the order satisfies $2\leq p\leq 7$ when $Y$ is either of $\textrm{Kum}_{n}$ type of of $OG6$ type, respectively $2\leq p\leq 23$ when $Y$ is either of $K3^{[n]}$ type or of $OG10$ type. Proposition 5.4 implies that the quotient $X:=Y/\varphi$ is log-Enriques of index $p$ when $p$ does not divide $n$, and in this case the second Betti number $b_{2}(X)$ coincides with the rank of the invariant sublattice $T_{\varphi}\subset H^{2}(Y,\mathbb{Z})$. As a consequence, this construction yields $2n$-dimensional log-Enriques varieties of index $p$ such that $(p,n)=1$ and with second Betti number $21\leq b_{2}(X)\leq b_{2}(Y)-$ 2, $b_{2}(X)=b_{2}(Y)\mod p-1$.
As observed in [12] and in [22], given a fixed component $Z\subset Y^{\varphi}$ of dimension $s$, the action of $\varphi$ linearizes near $q\in Z$ as:
$\left(\begin{matrix}I_{s}&&&&&&&\\ &\xi_{p}I_{s}&&&&&&\\ &&\xi_{p}^{a_{1}}&&&&&\\ &&&\xi_{p}^{p+1-a_{1}}&&&&\\ &&&&\ddots&&&\\ &&&&&\xi_{p}^{a_{t}}&&\\ &&&&&&\xi_{p}^{p+1-a_{t}}&\\ &&&&&&&\xi_{p}^{\frac{p+1}{2}}I_{2n-2s-2t}\end{matrix}\right)$
with $s\leq n$ by Lemma 5.5 and $1<a_{j}<p$ for all $1<a_{j}<p$; moreover, if $p=2$, $s=n$, hence $02n-2s-2t=$ 0.
By the Reid–Sheperd-Barron–Tai criterion for quotient singularities (see [51, Theorem 3.1] and [52, §4.11]), the quotient singularity in $\pi(q)$ is canonical (resp. terminal) if and only if the age $\alpha(\varphi)$ of $\varphi$ is $\geq 1$ (resp. $>1$). In this case, we compute
$\alpha(\varphi)=\left\{\begin{array}[]{cc}\frac{n}{2}&\ \mathrm{if}\ p=2\\ n-s+\frac{n}{p}&\ \mathrm{if}\ p>2\end{array}\right.$
@@
Indeed, if $s=n$, $\alpha(\varphi)=\frac{n}{p}$, but $p$ cannot be a divisor of $n$, hence $\alpha(\varphi)>1$ if $n>p$; otherwise, if $s<n$ then $p$ must be odd and $\alpha(\varphi)>n-s>1$. Canonical nonterminal singularities are never possible on these log-Enriques varieties, since $\alpha(\varphi)=1$ if and only if $p$ does not divide $n$, $s=n$ and $n=p$, impossible.
In the low dimensional case $n=2$, all the singularities of $X$ are terminal except for two-dimensional components of $X_{\rm sing}$, which are not even canonical.
6.2.2. Log-Enriques varieties simultaneously of IHS and CY type

### v2

@@ 6.1. Known examples of Enriques manifolds
There is a wide diversity of geometric constructions of Enriques surfaces, we refer for instance to [25, 28, 48]. However, in higher dimension quite few geometric constructions of nonsingular Enriques varieties exist. We briefly review the constructions given in [17, 51] and give some new examples using similar methods.
For any two-dimensional complex torus $A$ with origin $0\in A$, we denote by ${A}^{(n+1)}$ the $n+1$-th symmetric power of $A$, by $s\colon{A}^{(n+1)}\to A$ the summation map and by ${A}^{[n+1]}$ the Hilbert scheme of $n+1$ points on $A$. Consider the Hilbert–Chow morphism $\rho\colon{A}^{[n+1]}\to{A}^{(n+1)}$, the generalized Kummer variety (of dimension $2n$) of $A$ is the fiber ${\rm K}_{n}(A)\coloneqq(s\circ\rho)^{-1}(0)$.
@@
Let $E$ be an Enriques surface. Then for any $n\geq 2$, the Hilbert scheme ${E}^{[n]}$ of $n$ points on $E$ is an Enriques manifold of Calabi–Yau type, of dimension $2n$ and index two [51, Theorem 3.1].
@@
Let $A=E\times F$ be the product of two elliptic curves and consider the automorphism $\iota_{A}$ of $A$ defined by $\iota_{A}(x,y)=(-x+u,y+v)$. For any integer $n\geq 1$, the natural automorphism ${\iota_{A}}^{[n+1]}$ of ${A}^{[n+1]}$ respects the fiber ${\rm K}_{n}(A)$ if $u$ and $v$ are $(n+1)$-torsion points of $E$ and $F$ respectively. Moreover $\iota_{A}$ is an involution on $A$ if $v$ is a $2$-torsion point. Assuming thus that $n$ is odd, $u\in E[n+1]$ and $v\in F[2]$, it is easy to check that the natural automorphism ${\rm K}_{n}(\iota_{A})$ is a fixed point free involution on ${\rm K}_{n}(A)$ if $u$ is not an $\frac{n+1}{2}$-torsion point. Under these conditions, if $n$ is odd, by Propositions 5.1&5.4 the quotient ${\rm K}_{n}(A)/\langle{\rm K}_{n}(\iota_{A})\rangle$ is an Enriques manifold of symplectic type, of dimension $2n$ and index two.
@@ 6.1.3. Enriques involutions on K3 surfaces
Let $S$ be a K3 surface carrying an Enriques involution $\iota_{S}$, i.e. a fixed point free involution, and let $n\geq 1$ be an odd integer. Consider the natural involution ${\iota_{S}}^{[n]}$ induced on the Hilbert scheme ${S}^{[n]}$. Then, when $n$ is odd, the quotient ${S}^{[n]}/\langle{\iota_{S}}^{[n]}\rangle$ is an Enriques manifold of dimension $2n$, of symplectic type and index two. If $n$ is even, then by Proposition 5.4 the quotient has trivial canonical class so it is not an Enriques manifold. See Example 5.8 for a description of the case $n=2$.
@@
Let $S$ be a very general $K3$ surface carrying an Enriques involution $\iota_{S}$ as above By genericity we can assume that $\rho(S)=10$ and that the Picard lattice $\mathrm{Pic}(S)$ is $\iota_{S}$-invariant. Given a primitive Mukai vector $v=(r,l,\chi-r)\in H^{*}(S,\mathbb{Z})$ and a $v$-general polarization $H\in\mathrm{Pic}(S)$, the induced involution $\iota_{S}^{*}$ acts on the moduli space $M_{v,H}(S)$ of $H$-stable sheaves on $S$ with Mukai vector $v$. If $\chi$ is an odd integer, the quotient $M_{v,H}(S)/\langle\iota_{S}^{*}\rangle$ is an Enriques manifold of dimension $v^{2}+2$, of symplectic type and index two [51, Theorem 5.3]. This is a generalization of the previous construction.
@@ 6.1.5. Complete intersections of quadrics
For any even integer $n\geq 2$, consider the projective space $\mathbb{P}^{2n+1}$ with homogeneous coordinates $[x_{0}:\ldots:x_{n}:y_{0}:\ldots:y_{n}]$. For generic quadrics $Q_{j}(x),\widetilde{Q}_{j}(y)$ for $j=1,\ldots,n+1$, the complete intersection:
$Y\coloneqq\{[x:y]\,|\,Q_{j}(x)=\widetilde{Q}_{j}(y),\quad\forall j=1,\ldots,n+1\}$
is a nonsingular Calabi–Yau variety (we denote here $x:=(x_{0},\ldots,x_{n})$ and $y:=(y_{0},\ldots,y_{n})$). The involution $\iota$ of $\mathbb{P}^{2n+1}$ defined by $\iota(x_{i})=-x_{i}$ and $\iota(y_{i})=y_{i}$ for $i=0,\ldots,n$ fixes the two $n$-dimensional projective spaces defined by $\{x_{i}=0\,|\,i=0,\ldots,n\}$ and $\{y_{i}\,|\,i=0,\ldots,n\}$. For generic choices of the quadrics, this fixed locus does not meet $Y$ so the quotient $Y/\langle\iota\rangle$ is an Enriques manifold of Calabi–Yau type, of dimension $n$ and index two. Observe that if $n$ is odd, the same construction gives an involution preserving the volume form so the quotient is a Calabi–Yau manifold.
@@
Let $A=E\times E_{\omega}$ where $E$ is an elliptic curve, $\omega$ is a primitive $3$rd root of unity and $E_{\omega}=\frac{\mathbb{C}}{\mathbb{Z}\oplus\omega\mathbb{Z}}$. Take $u\in E$, $v\in E_{\omega}$ and consider the automorphism $f_{A}(x,y)=(x+u,\omega y+v)$. Assume that $u$ is a $3$-torsion point, so that $f_{A}$ has order $3$. Assuming that $u$ and $v$ are $(n+1)$-torsion points, the natural automorphism ${f_{A}}^{[n+1]}$ of ${A}^{[n+1]}$ respects the fiber ${\rm K}_{n}(A)$ and the necessary condition given by Proposition 5.1 imposes that $3$ divides $n+1$, so we write $n+1=3m$. If $m(2+\omega)v\neq 0$, the natural automorphism ${\rm K}_{n}(f_{A})$ is a fixed point free order three automorphism on ${\rm K}_{n}(A)$ and by Proposition 5.4, under these assumptions, if $3$ divides $n+1$ the quotient ${\rm K}_{n}(A)/\langle{\rm K}_{n}(f_{A})\rangle$ is an Enriques manifold of symplectic type, of dimension $2n$ and index three.
A variant of this construction, starting from a quotient $A=\frac{E\times E_{\omega}}{\mathbb{Z}/3\mathbb{Z}}$, produces for an appropriate choice of the automorphism a different family of Enriques manifolds of symplectic type as quotients of generalized Kummer manifolds, see [16, §4.3.7 - Type 6].
@@
Similarly as above, take $A=E\times E_{\mathrm{i}}$ where $E_{\mathrm{i}}=\frac{\mathbb{C}}{\mathbb{Z}\oplus\mathrm{i}\mathbb{Z}}$. Take $u\in E$, $v\in E_{\mathrm{i}}$ and consider the automorphism $f_{A}(x,y)=(x+u,\mathrm{i}y+v)$. Assume that $u$ is a $4$-torsion point, so that $f_{A}$ has order $4$. Assuming that $u$ and $v$ are $(n+1)$-torsion points, the natural automorphism ${f_{A}}^{[n+1]}$ of ${A}^{[n+1]}$ respects the fiber ${\rm K}_{n}(A)$ and the necessary condition given by Proposition 5.1 imposes that $4$ divides $n+1$, so we write $n+1=4m$. If $2mu\neq 0$ or if ${2m(1+\mathrm{i})v\neq 0}$, the natural automorphism ${\rm K}_{n}(f_{A})$ is a fixed point free order four automorphism on ${\rm K}_{n}(A)$ and by Proposition 5.4, under these assumptions, if $4$ divides $n+1$ the quotient ${\rm K}_{n}(A)/\langle{\rm K}_{n}(f_{A})\rangle$ is an Enriques manifold of symplectic type, of dimension $2n$ and index four.
@@ 6.2. Log-Enriques varieties of IHS type
By Proposition 5.4, any quotient of a $2n$-dimensional IHS manifold, with $n\geq 2$, by a purely non-symplectic automorphism is a log-Enriques variety as long as the order of the automorphism does not divide $n$. Here, we discuss some features of examples which can be constructed from some deformation families of IHS manifolds.
@@ 6.2.1. Log-Enriques varieties of prime index as quotients of IHS manifolds
Nonsymplectic automorphisms of prime order acting on IHS manifolds have now been classified for all known deformation families (see [14, 13, 16, 21, 22, 19] for $K3^{[n]}$ type, [47] for generalized Kummer manifolds, [36] for $OG_{6}$ manifolds and [12, 19] for $OG_{10}$ manifolds). Let $Y$ be an IHS manifold of dimension $2n\geq 4$ and let $\varphi\in\mathrm{Aut}(Y)$ be nonsymplectic of prime order $p$; in particular, the order satisfies $2\leq p\leq 7$ when $Y$ is either of $\textrm{Kum}_{n}$–type or of $OG_{6}$–type, respectively $2\leq p\leq 23$ when $Y$ is either of $K3^{[n]}$–type or of $OG_{10}$–type. Proposition 5.4 implies that the quotient $\pi\colon Y\to Y/\langle\varphi\rangle\eqqcolon X$ is log-Enriques of index $p$ when $p$ does not divide $n$, and in this case the second Betti number $b_{2}(X)$ coincides with the rank of the invariant sublattice $T_{\varphi}\subset H^{2}(Y,\mathbb{Z})$. As a consequence, this construction yields $2n$-dimensional log-Enriques varieties of index $p$ such that $(p,n)=1$ and with second Betti number $21\leq b_{2}(X)\leq b_{2}(Y)-$ 2, $b_{2}(X)=b_{2}(Y)\mod p-1$.
As explained in [14] and [24, Lemma 3.4(2)], given a fixed component $Z\subset Y^{\varphi}$ of dimension $s$, the action of $\varphi$ linearizes near a point $q\in Z$ as follows. If $p=2$ then $s=n$ since $Z$ is Lagrangian by [10, Lemma 1] and the action diagonalizes as:
$\rm{diag}\left(\underset{s\text{ times}}{\underbrace{1,\ldots,1}},\underset{s\text{ times}}{\underbrace{-1,\ldots,-1}}\right).$
If $p\geq 3$ then the diagonalization may be written as:
${\rm{diag}}\left(\underset{s\text{ times}}{\underbrace{{\underbracket{1,\xi_{p}},\ldots,\underbracket{1,\xi_{p}}}}},\underset{t\text{ times}}{\underbrace{{\underbracket{\xi_{p}^{a_{1}},\xi_{p}^{p+1-a_{1}}},\ldots,\underbracket{\xi_{p}^{a_{1}},\xi_{p}^{p+1-a_{t}}}}}},\underset{2n-2s-2t}{\underbrace{\xi_{p}^{\frac{p+1}{2}},\ldots,\xi_{p}^{\frac{p+1}{2}}}}\right),$
where $\xi_{p}$ is a primitive $p$-th root of unity, with $s\leq n$ by Lemma 5.5 since $Z$ is isotropic and $1<a_{j}<p$ for $1\leq j\leq t$. By the Reid–Sheperd-Barron–Tai criterion for quotient singularities (see [56, Theorem 3.1] and [57, §4.11]), the type of the quotient singularity at $\pi(q)$ is determined by its age (see [24]), denoted $\mathrm{age}(\varphi,\pi(q))$, as follows:
$\begin{cases}\text{klt}&\text{if }\mathrm{age}(\varphi,\pi(q))>0,\\ \text{canonical}&\text{if }\mathrm{age}(\varphi,\pi(q))\geq 1,\\ \text{terminal}&\text{if }\mathrm{age}(\varphi,\pi(q))>1,\\ \text{strictly klt}&\text{if }0<\mathrm{age}(\varphi,\pi(q))<1,\\ \text{Gorenstein}&\text{if }\mathrm{age}(\varphi,\pi(q))=1.\end{cases}$
We compute here:
$\mathrm{age}(\varphi,\pi(q))=\begin{cases}\frac{n}{2}&\text{if }p=2,\\ n-s+\frac{n}{p}&\text{if }p>2.\end{cases}$
@@
Indeed, if $s=n$, $\mathrm{age}(\varphi)=\frac{n}{p}$, but $p$ cannot be a divisor of $n$, hence $\mathrm{age}(\varphi)>1$ if $n>p$; otherwise, if $s<n$ then $p$ must be odd and $\mathrm{age}(\varphi)>n-s>1$. Canonical nonterminal singularities are never possible on these log-Enriques varieties, since $\mathrm{age}(\varphi)=1$ if and only if $p$ does not divide $n$, $s=n$ and $n=p$, this is impossible. In the low dimensional case $n=2$, all the singularities of $X$ are terminal except for two-dimensional components of $X_{\rm sing}$, which are not even canonical. Recall that a normal projective variety with numerically trivial canonical divisor is uniruled if and only if it has worse than canonical singularities see e.g. [27, Lemma 2.1] where this property is investigated for several log-Enriques varieties. We give in §6.2.2 an exemple of a log-Enriques variety of IHS type and index three which has strctly klt singularities, showing in this way that it is uniruled.
Remark 6.1.
By [61, Lemma 2.3] the maximum prime index for a log-Enriques surface is $19$. In higher dimension the prime index may be bigger. In fact in [13] it is shown that it exists and IHS manifold of $K3^{[2]}$-type admitting an automorphism of order $23$ acting purely nonsymplectically. By using the holomorphic Lefschetz formula one easily shows that the fixed locus is not empty and by Proposition 5.4 we get a log-Enriques variety of IHS-type and index $23$.
6.2.2. Log-Enriques varieties of IHS type from weak CY varieties